OpenStudy (anonymous):

You can buy DVDs at a local store fro $15.49 each. You can buy them at an online sore for $13.99 each plus $6 for shipping. How many DVDs can you buy for the same amount at the two stores?

OpenStudy (anonymous):

for this problem you need to create two equations; one that represents the amount of money it would cost to buy a certain amount of movies from the store, and another that shows the money it costs to buy a certain amount of movies online.

OpenStudy (anonymous):

d = 15.49 and o = 13.99 + 6?

OpenStudy (anonymous):

what you need in those equations is a variable to represent the number of movies

OpenStudy (anonymous):

m???

OpenStudy (anonymous):

because if I buy 2 movies from the store it will be 2 * $15.49

OpenStudy (anonymous):

so lets let x = the number of movies I want to buy and y is the amount of money it costs

OpenStudy (anonymous):

if I buy from the store I multiply the number of movies I want to buy times the price of each movie right?

OpenStudy (anonymous):

why not m and c?

OpenStudy (anonymous):

that works too. m for movies and c for cash?

OpenStudy (anonymous):

uhhuh ^^

OpenStudy (anonymous):

so the way we can represent the price of movies at the local store is with 15.49m = c

OpenStudy (anonymous):

that make sense?

OpenStudy (anonymous):

*nods*

OpenStudy (anonymous):

now we can make another equation that shows the price of buying movies online. Can you make that one?

OpenStudy (anonymous):

i can one sec

OpenStudy (anonymous):

13.99m + 6 = c

OpenStudy (anonymous):

good

OpenStudy (anonymous):

^^

OpenStudy (anonymous):

so now the problem asks us how many movies we can buy with the same amount of money at the two stores or for what value of m will both of our equations result in the same number c

OpenStudy (anonymous):

hmm can we combine them?

OpenStudy (anonymous):

and since 13.99m + 6 = c and 15.49m = c we can set these equations equal to each other because we're assuming c is the same number in both

OpenStudy (anonymous):

its like saying 2+2 = 4 and 1 +3 =4 so we can say that 1 + 3 = 2 + 2

OpenStudy (anonymous):

29.48m + 6 = c?

OpenStudy (anonymous):

oh ota, so 15.49m = 13.99m + 6

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

^^

OpenStudy (anonymous):

can you solve from here?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

let me know what you get

OpenStudy (anonymous):

m = 4

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

61.96 = c

OpenStudy (anonymous):

the question didnt ask for that, but yeah good job!

OpenStudy (anonymous):

it bothers me if i don't solve for all variables ... sorry